There are two trees in a garden, tree A and tree B and on both tress are some birds. The birds of tree A say to the birds of tree B that if one you comes to our tree, then our population will be the double of yours. Then the birds of tree B tell the birds of tree A that if one of you comes here then our poplation will be equal to that of yours. How many birds are in each tree?
Answer by ankor@dixie-net.com(15649)

There are two trees in a garden, tree A and tree B and on both trees are some birds.
The birds of tree A say to the birds of tree B that if one you comes to our tree, then our population will be the double of yours.
Then the birds of tree B tell the birds of tree A that if one of you comes here then our population will be equal to that of yours.
How many birds are in each tree?
:
let a = original number of birds in tree A
leg b = original number of birds in tree B
:
"The birds of tree A say to the birds of tree B that if one you comes to our tree, then our population will be the double of yours."
a + 1 = 2(b - 1)
a + 1 = 2b - 2
a = 2b - 2 - 1
a = 2b - 3
:
"Then the birds of tree B tell the birds of tree A that if one of you comes here then our population will be equal to that of yours."
a - 1 = b + 1
a = b + 1 + 1
a = b + 2
replace a with (2b-3), from the 1st statement
2b - 3 = b + 2
2b - b = 2 + 3
b = 5 birds on tree B, originally
then
a = 5 + 2
a = 7 birds on tree A
:
:
Check these solutions in the equation a = 2b - 3
7 = 2(5) - 3
7 = 10 - 3